A 1-mL solution of 0.05 M H2SO4, is diluted to 100 mL at 25 degrees C. What is the pH of the resulting solution?

a. 1

b. 2

c. 3

d. 4

Please explain the reasoning
Answer #1

PH would be 3 because

concentration of H ion will be from H2SO4 only

M1V1 = M2V2

0.05(1)=M2100

M2=510(4minus)

N= 2 (510(4minus)

N=110(3minus)

pH= -logH+

pH= -log110(3minus)

pH=3